OpenStudy (anonymous):

A circle has a diameter of 60 m. 1) The radius? 2) The circumference? 3)The area?

ILovePuppiesLol (ilovepuppieslol):

radius is half of the diameter

OpenStudy (luckyness):

Radius is half the diameter, so half of 60. 60/2=30 1)30

ILovePuppiesLol (ilovepuppieslol):

Circumference = π × diameter = 2 × π × radius. The circumference of a circle is the distance around it. The term is used when measuring physical objects, as well as when considering abstract geometric forms. When a circle's diameter is 1, its circumference is π.

OpenStudy (mathstudent55):

Use these formulas: \(d = 2r\) \(r = \dfrac{d}{2} \) \(C = \pi d = 2 \pi r\) \(A = \pi r^2\)

ILovePuppiesLol (ilovepuppieslol):

^ what i tried to say

OpenStudy (tgannon998):

Radius is 30 circumference is 2826